OpenStudy (anonymous):

add y/y^2-4y-5 +5/y+1

OpenStudy (anonymous):

You have to factor the first fraction's denominator.

OpenStudy (anonymous):

\[\frac{6 y-25}{y^2-4 y-5} \]

OpenStudy (anonymous):

\[= \frac{y}{(y-5)(y+1)} + \frac{5}{y+1} = \frac{y + 5(y-5)}{(y-5)(y+1)} = \frac{6y -25}{(y-5)(y+1)}\]